24 KN 3m- 4m

Solve using moment-area theorems.
Answer #1

all forms, La b [ey-27° atb+c= 3 -20=-1 C=1 So, a+b= 5/2 from bractical knowledge; a=1 & b=3/2 So, 0 k b / g 3 if H is double